With water from one hose, a swimming pool can be filled in 9 hours. A second, larger hose used alone can fill the pool in 4 hours. How long would it take to fill the pool if both hoses were used simultaneously?
Answer by checkley77(12844)

9*4/(9+4)=36/13=2.769 hours when both hoses are used.